### v1

@@ 4 Implementation
We implement the hSIDM scheme for efficiently describing DM self-interactions with arbitrary differential cross section in OpenGadget3 (see Groth et al., 2023, and the references therein), an upgraded version of the $N$-body code gadget-2 (Springel, 2005), following existing implementations for either purely forward-dominated (fSIDM) or large-angle scattering (rSIDM), respectively. In Sect. 4.1 we briefly review existing implementations, and then introduce the numerical algorithm underlying this work in Sect. 4.2. We discuss criteria for choosing the associated numerical time-steps in Sect. 4.3.
@@ 4.1 Existing implementations
In this section, we describe the numerical algorithm for the two existing methods reviewed in Sect. 2.3, starting with rSIDM for large-angle scattering and then turning to the purely forward-dominated fSIDM case.
4.1.1 Algorithm for rare large-angle scattering (rSIDM)
$N$-body simulations of DM self-interactions via large-angle scatterings, dubbed rSIDM, are well-established. Most studies assume the simplest case of isotropic scattering, but also an explicit sampling of deflection angles from a given distribution has been considered, see e.g. (Robertson et al., 2017b; Banerjee et al., 2020). Microscopic two-body interactions of DM particles are modelled by collisions between numerical $N$-body particles within rSIDM. Various approaches along these lines exist, differing in the computation of scattering probabilities (e.g. Koda & Shapiro, 2011; Vogelsberger et al., 2012; Rocha et al., 2013; Robertson et al., 2017b). Starting point is the interaction between two phase-space patches, represented by two numerical particles assumed to have equal mass. To simulate isotropic rSIDM, Fischer et al. (2021a) utilised the method presented by Rocha et al. (2013). The kernel function, $W$, characterises the density distribution of DM for a numerical particle in configuration space. Specifically, the scattering probability of a numerical particle $i$ to scatter with a particle $j$ having the mass $m_{j}$ can be expressed as
@@ 4.1.1 Algorithm for rare large-angle scattering (rSIDM)
where $\Delta\@vec{v}_{ij}$ is the relative velocity of the two numerical particles, $\Delta t$ denotes the time-step, and $\Lambda_{ij}=\int\mathrm{d}VW(|\@vec{x}-\@vec{x}_{i}|,h_{i})\,W(|\@vec{x}-\@vec{x}_{j}|,h_{j})$ gives the kernel overlap, with $h$ being the kernel size. Following Fischer et al. (2021a), we employ a spline kernel (Monaghan & Lattanzio, 1985) and choose $h$ adaptively such that it includes the $N_{\rm ngb}$ next neighbours. Based on the probability $P_{i}$, we determine whether two numerical particles scatter during the time-step. If a randomly selected number $x$ from the interval $[0,1]$ satisfies $x\leq P_{i}$, the particles scatter. Within the rSIDM scheme, the scattering angle is chosen randomly according to the differential cross section. For that purpose, it is convenient to consider the cumulative probability (as e.g. by Robertson et al., 2017b)
@@
4.1.2 Algorithm for frequent small-angle scattering (fSIDM)
The effective description of frequent, purely forward dominated scattering has been implemented by Fischer et al. (2021a). The self-interaction is described by a drag force and a diffusive random ‘kick’. As opposed to rSIDM, all pairs of numerical particles that are close enough to each other are affected by the interaction within a given time-step. Here ‘close enough’ means in practice that the kernel overlap is positive (i.e. $\Lambda_{ij}>0$). The individual ‘interaction’ of two numerical particles $i$ and $j$ within fSIDM is then composed of two steps. Firstly, as already discussed in Sect. 2.3, a drag force proportional to the (modified) transfer cross section is applied, with magnitude given by (see Eq. 9 by Fischer et al., 2021a) [^6] We note that the different normalisation of the cross section we use in this work, results in an additional factor of $1/2$ in Eq. (39) compared to Fischer et al. (2021a).
@@ 4.1.2 Algorithm for frequent small-angle scattering (fSIDM)
4.2 Implementation of the hybrid approach (hSIDM)
@@
In each time-step, for every pair with a positive kernel overlap $\Lambda_{ij}$, we simulate the small-angle scattering, before we model large scattering angles. Moreover, the pairwise computation of the particle interactions is done in a consecutive manner to conserve energy explicitly. This means that a numerical particle can not interact with multiple particles at the same time. Only after small- and large-angle scattering have been computed for the particles of a given pair, do further pairwise interactions with other particles follow. When the scatterings for all relevant pairs of numerical particles have been computed, the time-step is complete and the next one can follow.
@@
4.3 Choice of the time-step
Compared to conventional $N$-body simulations taking only gravity into account, the maximal possible size of a single time-step should be further constrained for SIDM to ensure accurate results, see for example Vogelsberger et al. (2012) and Fischer et al. (2021b, 2024b).
When simulating hSIDM and gravity, the upper limit on the time-step related to particle $i$ is determined as
@@ 4.3 Choice of the time-step
where $\Delta t_{{\rm grav.},i}$ is set by the gravitational acceleration following the criterion given by Springel (2005, see Eq. 34). The second one ($\Delta t_{{\rm large-angle},i}$) arises from the requirement that the scattering probability for large-angle scattering within a single time-step needs to be sufficiently smaller than unity, $P_{i}\ll 1$, [^8] This is the case even though our implementation can correctly describe multiple scatterings of a numerical particle per time-step. In Appendix G, we demonstrate and explain the required limitation of the scattering probability. leading to
@@
where $\kappa$ is a dimensionless accuracy parameter (we use $\kappa=0.01$), $m_{j}$ is the numerical particle mass, and $\Lambda_{ii}$ quantifies the maximum kernel overlap, determined from $\Lambda_{ij}$ with $j=i$. The third criterion ($\Delta t_{{\rm large-angle},i}$) arises from the effective treatment of frequent small-angle scattering. The description via a drag force and momentum diffusion requires that the typical distribution of deflection angles generated after a single time-step is still sufficiently narrow. In practice, this means that the contribution to the (modified) ‘transfer opacity’ (compare to Eq. (26)) from scatterings between $t$ and $t+\Delta t$, given by $\Delta\tau_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})=\rho v\Delta t\sigma_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})/m_{\chi}$ for a particle moving though a medium with mass density $\rho$ and relative velocity $v$, needs to be sufficiently small, $\Delta\tau_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})\ll{\cal O}(1)$. The origin of this criterion can be illustrated within the deflection test set-up discussed above, by considering the Molière approximation from Eq. (27), which corresponds to the result obtained within the effective small-angle approach. Here the transfer opacity determines the amount of Gaussian broadening. Hence, within hSIDM the condition $\Delta\tau_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})\ll{\cal O}(1)$ ensures that the contribution to the Gaussian width of the angular distribution from small-angle scatterings, and within a single time-step, is sufficiently small (compared to the full range of the azimuthal angle). This leads to
@@
where $\kappa$ is again an accuracy parameter (we use $\kappa=0.1$). This condition is analogous to the one considered in Fischer et al. (2024b) in the context of purely forward dominated scattering (i.e. fSIDM, for which $\sigma_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})$ is replaced by $\sigma_{\mathrm{\widetilde{T}}}$), see in particular Sect. 2.3 and Appendix B therein.
As already emphasised, the virtue of the hSIDM scheme is that for models with an enhancement of the differential cross section in the forward direction, the effective treatment of such scatterings is much more efficient as compared to the case when applying the sampling method (rSIDM approach) to the entire range of deflection angles. Indeed, within pure rSIDM, the time-step in Eq. (43) would contain $\sigma_{\rm tot}$, instead of $\sigma_{\mathrm{tot},>}(\theta_{\mathrm{c}})$ as for hSIDM. Due to $\sigma_{\rm tot}\gg\sigma_{\mathrm{tot},>}(\theta_{\mathrm{c}})$, this would lead to the requirement of excessively small time-steps in order to be able to treat also small-angle scatterings as individual collisions of numerical particles within a given time-step. Within hSIDM, only large-angle scatterings are treated in this way, while small-angle scatterings are treated separately by the effective drag force method. This relaxes the time-step requirement by one to two orders of magnitude for typical light mediator models.
@@ 5.1 Simulation set-up and basic validation
We consider the deflection test set-up discussed in Sect. 3.2, i.e. a beam of test particles moving through a target, and compare our numerical results to the analytical solution Eq. (18) for the time-evolution of the distribution function $f(t,\theta)$ of azimuthal angles of beam particles relative to the initial beam axis.
We set up an initial configuration with 8 00080008\,0008 000 test particles, initially moving with velocity $v=2\,\mathrm{km\,s}^{-1}$ in a common beam axis direction, through a medium described by 92 0009200092\,00092 000 target particles, with total mass $1\times 10^{10}\,\mathrm{M}_{\sun}$ in a cube with side length of $14\,\mathrm{kpc}$, corresponding to a background density $\rho=3.353\times 10^{6}\,\mathrm{M}_{\sun}\,\mathrm{kpc}^{-3}$. Here we use $N_{\mathrm{ngb}}=64$ neighbouring particles. Beam particles can scatter off target particles with a given differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$. We do not allow for scattering of beam with beam or target with target particles, and switch off gravity for the deflection test set-up. However, each beam particle can undergo multiple scatterings, with in principle any of the target particles.
@@
Since the parameters of the test set-up are chosen somewhat arbitrarily, we normalise all time-scales to a given reference time $t_{1}$, defined to be the dynamical time $t_{\rm dyn}$ on which the angular distribution is expected to evolve significantly, see Sect. 3.2. Concretely, we define $t_{1}$ such that the opacity $\tau(t)\equiv\rho vt\sigma/m_{\chi}$ satisfies $\tau(t_{1})=1$ for a generic cross section $\sigma/m_{\chi}=1\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. We note that this definition is insensitive to which type of cross section $\sigma$ stands for.
As a basic check of the implementation, we also consider an artificial case where each beam particle is allowed to scatter only once during the entire simulation. In this case the angular distribution simply follows the differential cross section, see Appendix D. After this check, we allow each beam particle to scatter multiple times, as determined according to the chosen scheme.
@@
In addition, we show the angular distribution obtained from the fSIDM scheme (c) within the no-recoil case (i) for three different times in Fig. 4 (solid lines). In this case, we compare to the Gaussian Molière distribution from Eq. (27) (dashed lines). Agreement between fSIDM and Eq. (27) is expected as long as the width of the Gaussian distribution is much smaller than $\uppi$, which is the case for the times $t$ shown in Fig. 4. As an additional check, the two panels in Fig. 4 show results for two values of the transfer cross section that differ by a factor $10$. Due to the absence of gravitational forces in the test set-up, the solutions for these two cases are expected to match when rescaling the simulation time by a factor $1/10$. We find that this is indeed the case, serving as an additional validation of the implementation of the time-step criterion. We note that this check is similar to the one performed by Fischer et al. (2021a), except that we consider the set-up without recoil (case (i)) here. This check validates that the effective treatment of small-angle scattering via a drag force and momentum diffusion leads to the expected Gaussian beam broadening. However, it does not provide a check of whether this treatment is a valid approximation for a given differential cross section and choice of critical angle. We turn to these checks next.
Figure 4: Validation of the implementation of angle-dependent SIDM based on the effective drag force and momentum diffusion approach (fSIDM scheme). Each panel shows the angular distribution of beam particles within the deflection test set-up at three different times $t$ (solid lines), compared to a Gaussian Molière distribution with width related to the transfer cross section (dashed lines), see Eq. (27). Here we assume beam-target scattering without recoil. The two panels correspond to two values of the transfer cross section, differing by a factor $10$. The distribution is expected to be identical at times rescaled by a factor $1/10$, which we observe to be the case.
@@ 5.2 Validity check of effective small-angle approach
The effective treatment of small-angle scattering via a drag force and momentum diffusion is valid provided the scattering is sufficiently forward dominated, and sufficiently frequent. In this sub-section we provide a quantitative check of these intuitive validity requirements. In order to expose the validity criteria most clearly, we consider the artificial differential cross section introduced in Eq. (4), for which all individual scatterings occur with a fixed scattering angle $\theta_{0}$ in the centre of mass frame. Details of this fixed scattering angle case are described in Appendix F. We solve the deflection test set-up for various fixed values of $\theta_{0}$. Since all scatterings occur with a single angle, we expect the effective small-angle approach to work the better the smaller $\theta_{0}$.
In order to check this expectation, we run two simulations for each value of $\theta_{0}$: one using the effective small-angle approach (fSIDM, scheme (b)), and one using the explicit sampling technique (rSIDM, scheme (c)). The latter is numerically more expensive, but yields the exact angular distribution function $f(t,\theta)$ at all times $t$ (within the numerical accuracy). This allows us to check whether the effective small-angle approach is valid, by comparing the angular distribution obtained from the fSIDM simulation to the rSIDM result.
@@
This finding is in line with the analytical discussion in Sect. 3.2.2, based on the analysis of the exact analytical solution of the deflection problem in the no-recoil case. Indeed, our numerical results precisely follow the expected behaviour, i.e. the fSIDM approach becomes valid only after some minimal time, that is required to lead to sufficiently many scatterings for the effective drag force description to apply. We find that the dependence of the minimal time $t_{\rm min}$ derived analytically in Eq. (32) on $\theta_{0}$ as well as on the overall size of the cross section (shown by solid lines in Fig. 5) agrees with our results for the numerically determined minimal time $t_{\rm min}(\theta_{0})$ (cross symbols). [^11] We note that the time $t_{\rm min}$ defined in Eq. (32) and the numerically determined $t_{\rm min}$ differ by a constant factor of order $10$ that is independent of $\theta_{0}$ and $\sigma_{\mathrm{T}}/m_{\chi}$. We attribute this to the precise definition for when the angular distributions obtained numerically within the fSIDM and rSIDM schemes agree with each other within numerical uncertainties. The solid lines in Fig. 5 are rescaled by this global common factor as compared to Eq. (32) in order to show that the dependence of the analytically derived and of the numerically extracted values of $t_{\rm min}$ on both $\theta_{0}$ and $\sigma_{\mathrm{T}}/m_{\chi}$ is (practically) identical. This demonstrates that the validity criteria derived analytically in Sect. 3.2.2 are applicable in practice. Moreover, our numerical results for the deflection test cases with and without recoil suggest that the validity criteria derived for the no-recoil set-up can be carried over to the case with recoil. This is an important observation, since all our later results are based on scatterings among identical particles, including recoil.
As can be observed in Fig. 5, the minimal time after which the fSIDM approach becomes applicable is the shorter the smaller $\theta_{0}$. This agrees with the intuitive notion that the fSIDM approach works the better the smaller the (typical) scattering angles are in individual collisions (which are given by $\theta_{0}$ for the differential cross section considered in this sub-section, and in general bounded from above by $\theta_{\mathrm{c}}$ within the hSIDM approach). As discussed in Sect. 3.2.2, the fSIDM approach is useful if $t_{\rm min}\ll t_{\rm dyn}$, where the latter is the dynamical time-scale over which the angular distribution evolves significantly. For the deflection set-up, $t_{\rm dyn}=(1,0.1,0.01)\cdot t_{1}$ for $\sigma_{\mathrm{T}}/m_{\chi}=(1,10,100)\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$, respectively. We find
@@
Altogether, this implies scatterings by angles $\theta\lesssim 1$ can be accurately treated within the effective small-angle (fSIDM) approach, since it becomes valid on time-scales that are one to two orders of magnitude shorter than the typical dynamical scale of the system.
Figure 5: Validity check of the effective small-angle approach (fSIDM, scheme b) as compared to an explicit treatment of angle-dependence (rSIDM, scheme c). Shown is the minimal time-scale $t_{\rm min}$ after which the angular distributions obtained for the deflection test problem solved with the fSIDM and the rSIDM method agree with each other. Cross symbols show the numerically determined minimal agreement time $t_{\rm min}$, for three values of $\sigma_{\mathrm{T}}/m_{\chi}$ and for the set-up with and without recoil in beam-target scattering, respectively. Lines show the analytical prediction for $t_{\rm min}$ from Eq. (32). Here we assume a differential cross section for which all individual scatterings occur with a fixed angle $\theta_{0}$ (shown on the $x$-axis). The numerically more efficient effective small-angle approach is advantageous if $t_{\rm min}$ is much smaller than the typical dynamical scale of the system, given by $t_{\rm dyn}=(1,0.1,0.01)\cdot t_{1}$ for $\sigma_{\mathrm{T}}/m_{\chi}=(1,10,100)\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$, respectively. This is safely the case for scatterings with $\theta\lesssim 1$. Details on the fixed-angle scattering are presented in Appendix F.

### v2

@@ 4 Implementation
We implement the hSIDM scheme for efficiently describing DM self-interactions with arbitrary differential cross section in OpenGadget3 (see Groth et al., 2023, and the references therein), an upgraded version of the $N$-body code gadget-2 (Springel, 2005), following existing implementations for either purely forward-dominated (fSIDM) or large-angle scattering (rSIDM), respectively. In Sect. 4.1 we briefly review existing implementations, and then introduce the numerical algorithm underlying this work in Sect. 4.2. We discuss criteria for choosing the associated numerical time steps in Sect. 4.3.
@@ 4.1 Existing implementations
In this section, we describe the numerical algorithm for the two existing methods reviewed in Sect. 2.3. We start with rSIDM for large-angle scattering and then turn to the purely forward-dominated fSIDM case.
4.1.1 Algorithm for rare large-angle scattering: rSIDM
$N$-body simulations of DM self-interactions via large-angle scatterings, dubbed rSIDM, are well-established. Most studies assume the simplest case of isotropic scattering, but also an explicit sampling of deflection angles from a given distribution has been considered (see e.g. (Robertson et al., 2017b; Banerjee et al., 2020)). Microscopic two-body interactions of DM particles are modelled by collisions between numerical $N$-body particles within rSIDM. Various approaches along these lines exist, differing in the computation of scattering probabilities (e.g. Koda & Shapiro, 2011; Vogelsberger et al., 2012; Rocha et al., 2013; Robertson et al., 2017b). Starting point is the interaction between two phase-space patches, represented by two numerical particles assumed to have equal mass. To simulate isotropic rSIDM, Fischer et al. (2021a) utilised the method presented by Rocha et al. (2013). The kernel function, $W$, characterises the density distribution of DM for a numerical particle in configuration space. Specifically, the scattering probability of a numerical particle $i$ to scatter with a particle $j$ having the mass $m_{j}$ can be expressed as
@@ 4.1.1 Algorithm for rare large-angle scattering: rSIDM
where $\Delta\@vec{v}_{ij}$ is the relative velocity of the two numerical particles, $\Delta t$ denotes the time step, and $\Lambda_{ij}=\int\mathrm{d}VW(|\@vec{x}-\@vec{x}_{i}|,h_{i})\,W(|\@vec{x}-\@vec{x}_{j}|,h_{j})$ gives the kernel overlap, with $h$ being the kernel size. Following Fischer et al. (2021a), we employ a spline kernel (Monaghan & Lattanzio, 1985) and choose $h$ adaptively such that it includes the $N_{\rm ngb}$ next neighbours. Based on the probability $P_{i}$, we determine whether two numerical particles scatter during the time step. If a randomly selected number $x$ from the interval $[0,1]$ satisfies $x\leq P_{i}$, the particles scatter. Within the rSIDM scheme, the scattering angle is chosen randomly according to the differential cross section. For that purpose, it is convenient to consider the cumulative probability (as e.g. by Robertson et al., 2017b)
@@
4.1.2 Algorithm for frequent small-angle scattering: fSIDM
The effective description of frequent, purely forward-dominated scattering has been implemented by Fischer et al. (2021a). The self-interaction is described by a drag force and a diffusive random ‘kick’. As opposed to rSIDM, all pairs of numerical particles that are close enough to each other are affected by the interaction within a given time step. Here ‘close enough’ means in practice that the kernel overlap is positive (i.e. $\Lambda_{ij}>0$). The individual ‘interaction’ of two numerical particles $i$ and $j$ within fSIDM is then composed of two steps. Firstly, as already discussed in Sect. 2.3, a drag force proportional to the (modified) transfer cross section is applied, with magnitude given by (see Eq. 9 by Fischer et al., 2021a) [^6] We note that the different normalisation of the cross section we use in this work, results in an additional factor of $1/2$ in Eq. (39) compared to Fischer et al. (2021a).
@@ 4.1.2 Algorithm for frequent small-angle scattering: fSIDM
4.2 Implementation of the hybrid approach: hSIDM
@@
In each time step, for every pair with a positive kernel overlap $\Lambda_{ij}$, we simulate the small-angle scattering, before we model large scattering angles. Moreover, the pairwise computation of the particle interactions is done in a consecutive manner to conserve energy explicitly. This means that a numerical particle cannot interact with multiple particles at the same time. Only after small- and large-angle scattering have been computed for the particles of a given pair, do further pairwise interactions with other particles follow. When the scatterings for all relevant pairs of numerical particles have been computed, the time step is complete and the next one can follow.
@@
4.3 Choice of the time step
Compared to conventional $N$-body simulations taking only gravity into account, the maximal possible size of a single time step should be further constrained for SIDM to ensure accurate results, see for example Vogelsberger et al. (2012) and Fischer et al. (2021b, 2024b). Next, we explain how we choose the time step.
When simulating hSIDM and gravity, the upper limit on the time step related to particle $i$ is determined as
@@ 4.3 Choice of the time step
where $\Delta t_{{\rm grav.},i}$ is set by the gravitational acceleration following the criterion given by Springel (2005, see Eq. 34). The second one ($\Delta t_{{\rm large-angle},i}$) arises from the requirement that the scattering probability for large-angle scattering within a single time step needs to be sufficiently smaller than unity, $P_{i}\ll 1$, [^8] This is the case even though our implementation can correctly describe multiple scatterings of a numerical particle per time step. In Appendix G, we demonstrate and explain the required limitation of the scattering probability. leading to
@@
where $\kappa$ is a dimensionless accuracy parameter (we use $\kappa=0.01$), $m_{j}$ is the numerical particle mass, and $\Lambda_{ii}$ quantifies the maximum kernel overlap, determined from $\Lambda_{ij}$ with $j=i$. The third criterion ($\Delta t_{{\rm large-angle},i}$) arises from the effective treatment of frequent small-angle scattering. The description via a drag force and momentum diffusion requires that the typical distribution of deflection angles generated after a single time step is still sufficiently narrow. In practice, this means that the contribution to the (modified) ‘transfer opacity’ (compare to Eq. (26)) from scatterings between $t$ and $t+\Delta t$, given by $\Delta\tau_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})=\rho v\Delta t\sigma_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})/m_{\chi}$ for a particle moving though a medium with mass density $\rho$ and relative velocity $v$, needs to be sufficiently small, $\Delta\tau_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})\ll{\cal O}(1)$. The origin of this criterion can be illustrated within the deflection test set-up discussed above, by considering the Molière approximation from Eq. (27), which corresponds to the result obtained within the effective small-angle approach. Here the transfer opacity determines the amount of Gaussian broadening. Hence, within hSIDM the condition $\Delta\tau_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})\ll{\cal O}(1)$ ensures that the contribution to the Gaussian width of the angular distribution from small-angle scatterings, and within a single time step, is sufficiently small (compared to the full range of the azimuthal angle). This leads to
@@
where $\kappa$ is again an accuracy parameter (we use $\kappa=0.1$). This condition is analogous to the one considered in Fischer et al. (2024b) in the context of purely forward-dominated scattering (i.e. fSIDM, for which $\sigma_{\mathrm{\widetilde{T}},<}(\theta_{\mathrm{c}})$ is replaced by $\sigma_{\mathrm{\widetilde{T}}}$), see in particular Sect. 2.3 and Appendix B therein.
As already emphasised, the virtue of the hSIDM scheme is that for models with an enhancement of the differential cross section in the forward direction, the effective treatment of such scatterings is much more efficient as compared to the case when applying the sampling method (rSIDM approach) to the entire range of deflection angles. Indeed, within pure rSIDM, the time step in Eq. (43) would contain $\sigma_{\rm tot}$, instead of $\sigma_{\mathrm{tot},>}(\theta_{\mathrm{c}})$ as for hSIDM. Due to $\sigma_{\rm tot}\gg\sigma_{\mathrm{tot},>}(\theta_{\mathrm{c}})$, this would lead to the requirement of excessively small time steps in order to be able to treat also small-angle scatterings as individual collisions of numerical particles within a given time step. Within hSIDM, only large-angle scatterings are treated in this way, while small-angle scatterings are treated separately by the effective drag force method. This relaxes the time step requirement by one to two orders of magnitude for typical light mediator models.
@@ 5.1 Simulation set-up and basic validation
We consider the deflection test set-up discussed in Sect. 3.2, consisting of a beam of test particles moving through a target. So we compare our numerical results to the analytical solution Eq. (18) for the time evolution of the distribution function $f(t,\theta)$ of azimuthal angles of beam particles relative to the initial beam axis.
We set up an initial configuration with 8 00080008\,0008 000 test particles, initially moving with velocity $v=2\,\mathrm{km\,s}^{-1}$ in a common beam axis direction, through a medium described by 92 0009200092\,00092 000 target particles, with total mass $1\times 10^{10}\,\mathrm{M}_{\sun}$ in a cube with side length of $14\,\mathrm{kpc}$, corresponding to a background density $\rho=3.353\times 10^{6}\,\mathrm{M}_{\sun}\,\mathrm{kpc}^{-3}$. Here we use $N_{\mathrm{ngb}}=64$ neighbouring particles. Beam particles can scatter off target particles with a given differential cross section $\mathrm{d}\sigma/\mathrm{d}\Omega$. We did not allow for scattering of the beam with beam or target with target particles, and we switched off gravity for the deflection test set-up. However, each beam particle can undergo multiple scatterings, with in principle any of the target particles.
@@
Since the parameters of the test set-up are chosen somewhat arbitrarily, we normalise all timescales to a given reference time $t_{1}$, defined to be the dynamical time $t_{\rm dyn}$ on which the angular distribution is expected to evolve significantly, see Sect. 3.2. Concretely, we define $t_{1}$ such that the opacity $\tau(t)\equiv\rho vt\sigma/m_{\chi}$ satisfies $\tau(t_{1})=1$ for a generic cross section $\sigma/m_{\chi}=1\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. We note that this definition is insensitive to which type of cross section $\sigma$ stands for.
As a basic check of the implementation, we also consider an artificial case where each beam particle is allowed to scatter only once during the entire simulation. In this case the angular distribution simply follows the differential cross section, see Appendix D. After this check, we allowed each beam particle to scatter multiple times, as determined according to the chosen scheme.
@@
In addition, we show the angular distribution obtained from the fSIDM scheme (c) within the no-recoil case (i) for three different times in Fig. 4 (solid lines). In this case, we compare to the Gaussian Molière distribution from Eq. (27) (dashed lines). Agreement between fSIDM and Eq. (27) is expected as long as the width of the Gaussian distribution is much smaller than $\uppi$, which is the case for the times $t$ shown in Fig. 4. As an additional check, the two panels in Fig. 4 show results for two values of the transfer cross section that differ by a factor of ten. Due to the absence of gravitational forces in the test set-up, the solutions for these two cases are expected to match when rescaling the simulation time by a factor of one-tenth. We find that this is indeed the case, serving as an additional validation of the implementation of the time step criterion. We note that this check is similar to the one performed by Fischer et al. (2021a), except that we consider the set-up without recoil (case (i)) here. This check validates that the effective treatment of small-angle scattering via a drag force and momentum diffusion leads to the expected Gaussian beam broadening. However, it does not provide a check of whether this treatment is a valid approximation for a given differential cross section and choice of critical angle. We turn to these checks next.
Figure 4: Validation of the implementation of angle-dependent SIDM based on the effective drag force and momentum diffusion approach (fSIDM scheme). Each panel shows the angular distribution of beam particles within the deflection test set-up at three different times $t$ (solid lines), compared to a Gaussian Molière distribution with width related to the transfer cross section (dashed lines), see Eq. (27). Here we assume beam-target scattering without recoil. The two panels correspond to two values of the transfer cross section, differing by a factor of ten. The distribution is expected to be identical at times rescaled by a factor of one-tenth, which we observe to be the case.
@@ 5.2 Validity check of effective small-angle approach
The effective treatment of small-angle scattering via a drag force and momentum diffusion is valid provided the scattering is sufficiently forward-dominated, and sufficiently frequent. In this sub-section we provide a quantitative check of these intuitive validity requirements. In order to expose the validity criteria most clearly, we consider the artificial differential cross section introduced in Eq. (4), for which all individual scatterings occur with a fixed scattering angle $\theta_{0}$ in the centre of mass frame. Details of this fixed scattering angle case are described in Appendix F. We solve the deflection test set-up for various fixed values of $\theta_{0}$. Since all scatterings occur with a single angle, we expect the effective small-angle approach to work the better the smaller $\theta_{0}$.
In order to check this expectation, we run two simulations for each value of $\theta_{0}$: one using the effective small-angle approach (fSIDM, scheme (b)), and one using the explicit sampling technique (rSIDM, scheme (c)). The latter is numerically more expensive, but yields the exact angular distribution function $f(t,\theta)$ at all times $t$ (within the numerical accuracy). This allowed us to check whether the effective small-angle approach is valid by comparing the angular distribution obtained from the fSIDM simulation to the rSIDM result.
@@
This finding is in line with the analytical discussion in Sect. 3.2.2, based on the analysis of the exact analytical solution of the deflection problem in the no-recoil case. Indeed, our numerical results precisely follow the expected behaviour; that is, the fSIDM approach becomes valid only after some minimal time that is required to lead to sufficiently many scatterings for the effective drag force description to apply. We find that the dependence of the minimal time $t_{\rm min}$ derived analytically in Eq. (32) on $\theta_{0}$ as well as on the overall size of the cross section (shown by solid lines in Fig. 5) agrees with our results for the numerically determined minimal time $t_{\rm min}(\theta_{0})$ (cross symbols). [^11] We note that the time $t_{\rm min}$ defined in Eq. (32) and the numerically determined $t_{\rm min}$ differ by a constant factor of the order of ten, that is, independent of $\theta_{0}$ and $\sigma_{\mathrm{T}}/m_{\chi}$. We attribute this to the precise definition for when the angular distributions obtained numerically within the fSIDM and rSIDM schemes agree with each other within numerical uncertainties. The solid lines in Fig. 5 are rescaled by this global common factor as compared to Eq. (32) in order to show that the dependence of the analytically derived and of the numerically extracted values of $t_{\rm min}$ on both $\theta_{0}$ and $\sigma_{\mathrm{T}}/m_{\chi}$ is (practically) identical. This demonstrates that the validity criteria derived analytically in Sect. 3.2.2 are applicable in practice. Moreover, our numerical results for the deflection test cases with and without recoil suggest that the validity criteria derived for the no-recoil set-up can be carried over to the case with recoil. This is an important observation, since all our later results are based on scatterings among identical particles, including recoil.
As can be observed in Fig. 5, the minimal time after which the fSIDM approach becomes applicable is the shorter the smaller $\theta_{0}$. This agrees with the intuitive notion that the fSIDM approach works the better the smaller the (typical) scattering angles are in individual collisions (which are given by $\theta_{0}$ for the differential cross section considered in this sub-section, and in general bounded from above by $\theta_{\mathrm{c}}$ within the hSIDM approach). As discussed in Sect. 3.2.2, the fSIDM approach is useful if $t_{\rm min}\ll t_{\rm dyn}$, where the latter is the dynamical timescale over which the angular distribution evolves significantly. For the deflection set-up, $t_{\rm dyn}=(1,0.1,0.01)\cdot t_{1}$ for $\sigma_{\mathrm{T}}/m_{\chi}=(1,10,100)\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$, respectively. We find
@@
Altogether, this implies scatterings by angles $\theta\lesssim 1$ can be accurately treated within the effective small-angle (fSIDM) approach. This is valid on timescales that are one to two orders of magnitude shorter than the typical dynamical scale of the system.
Figure 5: Validity check of the effective small-angle approach (fSIDM, scheme b) as compared to an explicit treatment of angle dependence (rSIDM, scheme c). Shown is the minimal timescale $t_{\rm min}$ after which the angular distributions obtained for the deflection test problem solved with the fSIDM and the rSIDM method agree with each other. Cross symbols show the numerically determined minimal agreement time $t_{\rm min}$, for three values of $\sigma_{\mathrm{T}}/m_{\chi}$ and for the set-up with and without recoil in beam-target scattering, respectively. Lines show the analytical prediction for $t_{\rm min}$ from Eq. (32). Here we assume a differential cross section for which all individual scatterings occur with a fixed angle $\theta_{0}$ (shown on the $x$-axis). The numerically more efficient effective small-angle approach is advantageous if $t_{\rm min}$ is much smaller than the typical dynamical scale of the system, given by $t_{\rm dyn}=(1,0.1,0.01)\cdot t_{1}$ for $\sigma_{\mathrm{T}}/m_{\chi}=(1,10,100)\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$, respectively. This is safely the case for scatterings with $\theta\lesssim 1$. Details on the fixed-angle scattering are presented in Appendix F.